OpenStudy (anonymous):

WILL GIVE MEDAL!!! Explain how to find instantaneous velocity and find it for s(t) = -3 + 4t^2 when t = 5 seconds. @Mertsj @zepdrix @Hero @mathstudent55

OpenStudy (anonymous):

http://www.wikihow.com/Calculate-Instantaneous-Velocity Go to website above to find out how to find instantaneous velocity

OpenStudy (anonymous):

@raesal

hero (hero):

@raesal Just evaluate s(5).

OpenStudy (anonymous):

So -3+4(5)^2=97. The instantaneous velocity would be 97? @Hero

hero (hero):

Actually, you should find v(t) first, then evaluate v(5).

OpenStudy (anonymous):

How do I find v(t)? @Hero

hero (hero):

v(t) = s'(t)

OpenStudy (anonymous):

so v(t)= -3 + 4t^2?

hero (hero):

You have to calculate s'(t) first. In other words, take the derivative of s(t) to get v(t).

OpenStudy (anonymous):

The derivative of -3 + 4t^2 is 8t. so 8t=vt? Then what?

hero (hero):

Now evaluate v(5).

OpenStudy (anonymous):

so 8*5=40. so the instantaneous velocity is 40?

OpenStudy (anonymous):

But I have the correct number, 40, correct? @Hero